OpenStudy (anonymous):

Multivariable Calculus: Arc length parametrization of a curve; question in reply

OpenStudy (anonymous):

Consider the curve parameterized by \[r(t) = (e^{t}, \sqrt{2}t, e^{-t}), t \ge 0\] Find an arc length parametrization of the curve.

OpenStudy (anonymous):

I've been able to find the arc length, \[s = e^{t} - e^{-t}\] I hope this is right. Anyways, I believe I have to solve for t now, but I'm really stuck on this part. Any hints would be great, thanks!

OpenStudy (anonymous):

Find R'(T) then find |r'(t)| arc length =\[\int\limits_{a}^{b} |(r't)| dt\]

OpenStudy (anonymous):

use this formula when a curve is given in parametric form

OpenStudy (anonymous):

kay, thanks I'l try that!

OpenStudy (dape):

That's the wrong arc length parametrization.

OpenStudy (anonymous):

well i've already found the arc length or was I wrong?

OpenStudy (anonymous):

Oh okay

OpenStudy (dape):

From the given r(t) we get \[ r'(t)=(e^t,\sqrt{2},-e^{-t}) \] The arc length parametrization is now: \[ s(t)=\int{||r'(t)||dt} \]

OpenStudy (anonymous):

I'm still getting

OpenStudy (anonymous):

\[s(t) = e^{t} - e^{-t}\] so is that the arc length parametrization?

OpenStudy (dape):

Ah, sorry, yeah, that's right, I used hyperbolic functions when I integrated so my answer looked different.

OpenStudy (dape):

I got \( s(t)=2\sinh t \).

OpenStudy (dape):

Which is the same

OpenStudy (anonymous):

Okay! Thanks! So would that be my answer? I thought that the arc length par'n was found by plugging the inverse of s(t) into r. So I'd have to solve for t, if that makes sense.

OpenStudy (dape):

You're right, s(t) of course gives the arc length for a given t. And we want to put in a length and get out the corresponding point at that arc length of the curve. The inverse is just t(s) = arcsinh(s/2), just plug that into the original parametrization. I think I should go get some sleep, it's 4 am here. Good luck!

OpenStudy (anonymous):

Oh wow, thanks for helping when it's so 4am where you are! Goodnight and thanks again

OpenStudy (dape):

No problem, you basically solved it without any help anyways =)